# GeoBenchX: Benchmarking LLMs for Multistep Geospatial Tasks

### v1

@@ 5.1 Matching-based evaluator
The 2 widely used evaluation approaches are execution-based and matching-based. However, comparing directly the outputs of the agent on geospatial analytical tasks would be quite challenging because the outputs contain multiple images/maps, tables, numbers, a narrative and sometime all of it as in the section 4.2 example. Annotation of such ground truth is not practical in this case, so we opted for matching-based approach. We used LLM-as-Judge with multiple LLM judges to compare the solutions generated by the agent (“candidate solutions”) and reference solutions for the same task. To leverage the LLMs ability to analyze code, we represented both reference and candidate solutions as snippets of Python code as in 4.2. We asked LLM judges to compare semantic equivalence of the solution; this took care of situations where reference and candidate solutions had different order of steps, different arguments for tool-calls or even took a different path to the same analytical outcome.
@@
We also compared the lengths of the reference and candidate solutions, since the agent often takes some extra steps to figure out a path to answering the user question, and such extra steps are often costly.
@@
For evaluation of the solutions, we used a panel of judges: Sonnet 3.5, GPT-4o, and Gemini 2.0. The evaluator prompt included: persona setting, description of all tools, scoring taxonomy, instructions on how evaluate solutions, examples of evaluated solutions, corresponding scores, and reasoning behind each score. The evaluator was instructed to analyze the candidate solution for the task and produce (1) reasoning for matching any of the reference solutions for the task and the candidate solution and (2) the matching score (‘2’: match, ‘0’: no match, ‘1’: partial match)
5.2 Evaluating the evaluator
@@
We evaluated and tuned the evaluator itself on a set of 44 tasks randomly sampled with stratification of all tasks, respecting task groupings. We manually scored candidate solutions for these tasks, providing a matching score and reasoning for it. We then ran the evaluator on the same tasks candidate solutions and compared human annotator match scores, and the evaluator produced match scores.
Among the three models used as judges, the Sonnet 3.5 model matching scores were closest to human annotator scores. scoring (0.91 of the scores corresponded to manual scores), followed by GPT-4o (0.86) and Gemini 2.0 (0.82).
@@
In the below results section, we use all the 3 above LLMs as judges and average their match frequency estimates.
6 LLMs evaluation results
@@
The following models were tested in the solver agent: gpt-4o-2024-08-06, gpt-4o-mini-2024-07-18, o3-mini-2025-01-31, gemini-2.0-flas%h-001, claude-3-5-sonnet-20241022, claude-3-5-haiku-20241022, claude-3-7-sonne%t-20250219. Each LLM was tested on the same set of tasks, and all models were tested at temperature 0 (except o3-mini, which does not have a temperature parameter).
The frequency of each of the match scores was calculated (1) separately for the solvable tasks (104 tasks) and the unsolvable tasks (98 tasks) (2) for each group of tasks according to 4.
To mark unsolvable tasks, the agent was provided with a rejecttask tool that it could call either upfront or after some steps if it reached the conclusion that the task cannot be solved with available datasets and tools.
@@
To benchmark how efficient an LLM is at solving the task compared to the reference solution, we calculated the mean difference between the length (number of steps) of the reference solution (or average length of reference solutions if multiple) and the length of candidate solution. We express this difference relative to the length of the reference solution (so 0.5 means 50% more steps in the candidate solution than the reference).
The first observation is that models that are more successful in identifying unsolvable tasks tend to be less successful in dealing with solvable tasks and vice versa (Table 1, Table 2). Anthropic models perform better on solvable tasks, but prioritize giving any answer at all over rejecting tasks, Sonnet 3.5 is the most balanced from the three in this aspect. Gemini 2.0 performs similarly on both kinds of tasks. Although the three OpenAI models are better at rejecting tasks than solving them, among them, the GPT-4o is well balanced. Overall, on solvable tasks, Sonnet 3.5 turned out to be the best model, followed by Sonnet 3.7, GPT-4o, Gemini 2.0 and Haiku 3.5 (Table 2). We can also see that for unsolvable tasks, the candidate solutions are much longer, indicating LLMs trying to find a solution and eventually giving up. For solvable tasks, candidate solutions are about the same length as reference ones, with notable exceptions of Sonnet 3.7 that tends to produce longer solutions than needed and o3-mini that tends to reject tasks more frequently than other models. The detailed statistics for each group of tasks are in tables 3, 4, 5, 6.
Model No Partial Match $\Delta$ in
match match solution
length
Sonnet 3.7 0.84 0.02 0.14 6.47
Sonnet 3.5 0.61 0.02 0.37 2.51
Haiku 3.5 0.79 0.01 0.20 4.39
GPT-4o 0.38 0.03 0.59 2.45
GPT-4o mini 0.37 0.02 0.62 2.71
o3-mini 0.23 0.02 0.76 0.20
Gemini 2.0 0.54 - 0.46 1.83
Table 1: Success rate in identifying the unsolvable tasks (98 tasks from total of 202)
Model No Partial Match $\Delta$ in
match match solution
length
Sonnet 3.7 0.27 0.23 0.50 0.39
Sonnet 3.5 0.28 0.19 0.53 -0.05
Haiku 3.5 0.34 0.23 0.43 0.03
GPT-4o 0.41 0.15 0.44 -0.01
GPT-4o mini 0.70 0.13 0.16 -0.03
o3-mini 0.69 0.10 0.21 -0.59
Gemini 2.0 0.43 0.14 0.43 -0.03
Table 2: Success rate in solvable tasks (104 tasks from total of 202)
Model No Partial Match $\Delta$ in
match match solutions
length
Sonnet 3.7 0.23 0.11 0.66 0.91
Sonnet 3.5 0.17 0.15 0.72 0.29
Haiku 3.5 0.25 0.21 0.54 0.51
GPT-4o 0.24 0.13 0.63 0.31
GPT-4o mini 0.49 0.19 0.32 0.31
o3-mini 0.50 0.14 0.36 -0.39
Gemini 2.0 0.25 0.06 0.68 0.16
Table 3: Success rate in solving ‘Merge-visualize’ tasks (36 tasks from total of 202)
Model No Partial Match $\Delta$ in
match match solution
length
Sonnet 3.7 0.50 0.14 0.36 2.37
Sonnet 3.5 0.30 0.10 0.60 0.40
Haiku 3.5 0.59 0.09 0.33 2.11
GPT-4o 0.26 0.09 0.65 0.29
GPT-4o mini 0.54 0.07 0.39 0.60
o3-mini 0.48 0.05 0.46 -0.38
Gemini 2.0 0.35 0.12 0.53 0.17
@@
Table 4: Success rate in solving ‘Process-merge-visualize’ tasks (56 tasks from total of 202)
Model No Partial Match $\Delta$ in
match match solution
length
Sonnet 3.7 0.55 0.14 0.31 3.42
Sonnet 3.5 0.48 0.12 0.40 1.73
Haiku 3.5 0.50 0.17 0.33 2.49
GPT-4o 0.50 0.08 0.42 1.95
GPT-4o mini 0.69 0.05 0.26 2.44
o3-mini 0.55 0.07 0.39 -0.3
Gemini 2.0 0.63 0.07 0.30 2.07
Table 5: Success rate in solving ‘Spatial operations’ tasks (53 tasks from total of 202)
Model No Partial Match $\Delta$ in
match match solution
length
Sonnet 3.7 0.81 0.13 0.06 6.05
Sonnet 3.5 0.76 0.10 0.13 2.15
Haiku 3.5 0.82 0.06 0.12 3.06
GPT-4o 0.56 0.08 0.36 2.00
GPT-4o mini 0.44 0.03 0.53 1.62
o3-mini 0.39 - 0.61 0.17
Gemini 2.0 0.64 0.04 0.32 0.91
@@
Table 6: Success rate in solving ‘Heatmaps, contour lines’ tasks (54 tasks from total of 202)
@@
Sonnet 3.7 24 269
Sonnet 3.5 12 157
Haiku 3.5 15 203
GPT-4o 7.5 57
GPT-4o mini 8 55
o3-mini 3.5 561
Gemini 2.0 7 40
Table 7: Tokens used for generation of solutions for the whole set
6.1 Error Analysis
@@
(1) Poor understanding of geometrical relations: mistaking centroids for full polygons, attempting to get centroids of lines, and not understanding that centroids can be outside object boundaries, etc, etc. (2) Outdated world knowledge – good for established international groupings but fails with recent changes (e.g., BRICS expansion in 2024) or ambiguous groupings (Upper/Lower Nile countries). (3) Over-reliance on world knowledge - using internally known spelling of names rather than dataset values for filtering, despite having tools to check column values. (4) Inefficient data manipulation – merging without pre-filtering (e.g., merging with all USA counties instead of first filtering to specific state). (5) Skipping geographic filtering steps – keeping continent-level data when only country-level was needed. (6) Attempting to solve with partial data, trying to be helpful by using whatever geography was available instead of rejecting tasks with insufficient data coverage. (7) Modifying tasks to make them solvable rather than solving the original problem.
7 Conclusions

### v2

@@ 5.1. Matching-based evaluator
To leverage the LLMs ability to comprehend code (Nikiema et al., 2025; Haroon et al., 2025; Ma et al., 2024), we represented both reference and candidate solutions as snippets of Python code as in 4.3. In code evaluation, the two widely used approaches are: execution-based and matching-based, both used for the evaluation of LLM-generated code (Wang and Chen, 2023; Anjum Haque et al., 2023). Since the outputs of agent on geospatial analytical tasks contain multiple images/maps, numbers, a narrative, and sometimes all of it as in the section 4.3 example, reliably comparing them via execution-based approach would be challenging and annotation of tasks with such ground truth is rarely practical. We opt for the matching-based approach for practicality and efficiency. We used reference-based LLM-as-Judge with multiple LLM judges to compare the solutions generated by the agent (“candidate solutions”) and reference solutions for the same task. Compared to human evaluation, the use of LLMs to compare the generated solution and reference is cost- and time-efficient (Chiang and Lee, 2023); if multiple rounds of evaluation are needed due to changes in tasks, their number, or evaluated LLMs, it offers reproducibility (Chiang and Lee, 2023) and scalability (Zheng et al., 2023).
We asked LLM judge to compare semantic equivalence of the solution; this took care of situations where reference and candidate solutions had different order of steps, different arguments for tool-calls or even took a different path to the same analytical outcome.
@@
We also compared the lengths of the reference and candidate solutions, since the agent often takes some steps to figure out a path to answering the user question, and such extra steps might be costly (see 6.4). Looking into solution code also provides interesting insights into how LLM solves the tasks and allows agent tuning (which we did not do for this benchmarking).
@@
For evaluation of the solutions we used a panel of judges: Claude Sonnet 3.5, GPT-4.1 and Gemini 2.5 Pro Preview. The evaluator prompt included: persona setting, description of all tools, scoring taxonomy, instructions on how evaluate solutions, examples of evaluated solutions, corresponding scores and reasoning behind each score. The evaluator was instructed to analyze the candidate solution for the task and produce (1) reasoning for matching any of the reference solutions for the task and the candidate solution and (2) the matching score (’2’: match, ’0’: not a match, ’1’: partial match).
5.2. Evaluating the evaluator
@@
We evaluated and tuned the evaluator itself on the set of 50 tasks (about 25% of the benchmark set) randomly sampled with stratification from all tasks, respecting the task groups. We manually scored candidate solutions for these tasks, providing a match score and reasoning for it. We then ran the evaluator on the same tasks candidate solutions and we compared human annotator match scores and evaluator produced match scores.
Nine models were tested as judges (Gemini 2.0 Flash, Gemini 2.5 Pro Preview, Gemini 2.5 Flash Preview; Claude Sonnet 3.5, 3.7, 4; GPT-4o, GPT-4.1, GPT-4.1 mini) and from each family of models the one with the results closest to human annotator scores was selected for the panel: GPT-4.1 with 96% of scores matching the manual scores, Claude Sonnet 3.5 (94%) and Gemini 2.5 Pro Preview (88%) (see Figure 1). The top models of each family were selected rather than simply the top three models to reduce variance in the evaluation results due to differences in training and tuning of the models by different providers.
Figure 1. Testing of LLMs as judges:
@@
In the evaluator, we use all the 3 above LLMs as judges and average their match frequency estimates.
6. LLMs evaluation results
6.1. Experiment procedure
@@
Nine models were tested in the solver agent: gemini-2.0-flash-001, gemini-2.5-pro-preview-05-06, claude-3-5-haiku-20241022, claude-3-5-sonnet-20241022, claude-sonnet-4-20250514, gpt-4o-2024-08-06, gpt-4.1-2025-04-14, o3-mini-2025-01-31, o4-mini-2025-04-16. Each LLM was tested on the same benchmark set of tasks, at temperature 0 (except o3-mini, o4-mini which do not support the temperature parameter).
The frequency of each of the match scores was calculated:
  1. (1)
    separately by each group of tasks as per 4.1 and
  2. (2)
    for solvable tasks (123 tasks) and unsolvable tasks (79 tasks) across all groups.
@@
To benchmark how efficient an LLM is at solving a task compared to the reference solution, we also calculated the mean difference between the length (number of steps) of the reference solution for the task (or average length of reference solutions if multiple) and the candidate solution. We express this difference relative to the length of candidate solution (so 0.5 means 50% more steps in the candidate solution).
The o3-mini model generated empty solutions (no tool calls, just thinking) for 103 tasks and rejected 75 tasks, and while the accuracy in identification of unsolvable tasks thus was 0.75, the model was excluded from further analysis due to obvious inefficiency in using tools to solve tasks.
6.2. Performance on solvable and unsolvable tasks
The first observation is that some models demonstrate a trade-off between the ability to identify and reject unsolvable tasks and the ability to solve tasks (Table 1, Table 2, Table 7, Figure 2). Claude Sonnet 4 and Haiku 3.5 perform better on solvable tasks but prioritize giving any answer over rejecting tasks, with Claude Sonnet 4 being a prominent case with a success rate on solvable tasks more than three times higher than on unsolvable ones. Gemini 2.5 Pro Preview is somewhat better at detecting unsolvable tasks than solving solvable ones. GPT-4o and o4-mini are substantially better in detecting unsolvable tasks than in solving feasible ones. In each family of models, there is one that demonstrated similar accuracy on both types of tasks in the experiment: Claude Sonnet 3.5, Gemini 2.0 Flash and GPT-4.1.
We can also see, that for unsolvable tasks, the candidate solutions are much longer indicating LLMs trying to find a solution and eventually giving up. For solvable tasks, candidate solutions are about the same length as reference ones, with the exception of Claude Sonnet 4 which tends to produce longer solutions than needed.
Overall, in solvable tasks, Claude Sonnet 4 turned out to be the best model, followed by o4-mini, Claude Sonnet 3.5, GPT-4.1, and the rest of the models were very close in their success rates (Table 2). However, the highest success rate is only slightly more than half (0.55 and 0.51), and most of the tested models are successful in less than half of the cases (note again that we specifically chose a trivial agent architecture to benchmark the ability of LLMs themselves rather than a particular agentic architecture). Some models have a relatively high proportion of solutions that partially match ground truth solutions (0.16-0.22 for Claude Haiku 3.5, GPT-4.1, Claude Sonnet 4, and Gemini 2.5 Pro Preview), meaning these solutions are very close to correct and models success rates might be increased with minor changes to the agent and tools.
Table 1. Success rate in identifying the unsolvable tasks
(79 tasks from total of 202) Model No Partial Match $\Delta$ in match match solution length Gemini 2.0 Flash 0.60 0.00 0.40 2.3 Gemini 2.5 Pro Preview 0.45 0.00 0.55 2.0 Claude Haiku 3.5 0.78 0.00 0.22 4.2 Claude Sonnet 3.5 0.53 0.00 0.47 2.5 Claude Sonnet 4 0.83 0.00 0.17 7.6 GPT-4o 0.38 0.00 0.62 2.5 GPT-4.1 0.57 0.00 0.43 4.5 o3-mini 0.25 0.00 0.75 0.0 o4-mini 0.10 0.00 0.90 0.5
Table 2. Success rate in solvable tasks
(123 tasks from total of 202) Model No Partial Match $\Delta$ in match match solution length Gemini 2.0 Flash 0.52 0.08 0.39 0.0 Gemini 2.5 Pro Preview 0.44 0.16 0.39 0.0 Claude Haiku 3.5 0.38 0.22 0.40 0.2 Claude Sonnet 3.5 0.41 0.12 0.47 -0.1 Claude Sonnet 4 0.27 0.18 0.55 0.7 GPT-4o 0.51 0.12 0.38 0.1 GPT-4.1 0.38 0.20 0.42 0.4 o4-mini 0.38 0.11 0.51 0.1
Figure 2. Success rate (share of matches with reference solutions) across LLMs and task types:
6.3. Performance by task groups
We then examine LLMs performance by task groups without distinguishing between solvable and unsolvable tasks within each group. In the ’Merge - visualize’ task group (Table 3), the easiest and, mostly, solvable, almost all models achieved accuracy over 0.6. The o4-mini had the highest accuracy, followed by Claude Sonnet 3.5, 4, all with success rates over 0.7.
Table 3. Success rate in solving ’Merge - visualize’ tasks
(36 tasks from total of 202) Model No Partial Match $\Delta$ in match match solution length Gemini 2.0 Flash 0.32 0.00 0.68 0.2 Gemini 2.5 Pro Preview 0.20 0.13 0.67 0.7 Claude Haiku 3.5 0.25 0.15 0.60 0.6 Claude Sonnet 3.5 0.17 0.11 0.72 0.3 Claude Sonnet 4 0.19 0.09 0.71 1.3 GPT-4o 0.32 0.08 0.60 0.4 GPT-4.1 0.28 0.12 0.60 0.8 o4-mini 0.20 0.03 0.77 0.1
In the ’Process - merge - visualize’ task group, accuracy starts to decrease while the solution lengths increase (Table 4). The top three models by accuracy achieve success rates over 0.6: GPT-4o, Claude Sonnet 3.5, and o4-mini, with o4-mini notably outperforming the other eight models at 0.75 accuracy. The rest have accuracy 0.5 or lower.
In this task group, the tendency of Claude Sonnet 4 to produce lengthy tasks becomes apparent: its solutions on average are 330% longer than the reference solutions. Claude Haiku 3.5, the next model by mean difference in length with reference solutions, has the solutions 190% longer on average. The longer solutions usually mean repeated or failed steps, unnecessary or failed visualizations or unnecessary data loads early in the solution. This is associated with more errors, with the final response sometimes being unusable even if it contains the answer, or with inability to complete solution within the recursion limit. It also confuses evaluators that tend to penalize excessively long solutions despite the instructions not to. Most common problems for these two models were attempts to provide solutions for unsolvable tasks, errors in data filtering, selection of datasets and dates for mapping, lack of awareness of recent updates (like expansion of BRICS) even for models released after the events.
@@ 6.3. Performance by task groups
Table 4. Success rate in solving ’Process - merge - visualize’ tasks
(56 tasks from total of 202) Model No Partial Match $\Delta$ in match match solution length Gemini 2.0 Flash 0.46 0.06 0.48 0.04 Gemini 2.5 Pro Preview 0.45 0.05 0.50 1.0 Claude Haiku 3.5 0.54 0.09 0.36 1.9 Claude Sonnet 3.5 0.36 0.00 0.61 0.3 Claude Sonnet 4 0.50 0.08 0.41 3.3 GPT-4o 0.32 0.07 0.61 0.2 GPT-4.1 0.44 0.06 0.51 1.3 o4-mini 0.23 0.03 0.75 0.4
In the third task group, ’Spatial operations’ only o4-mini achieves a success rate 0.56 (Table 5). However, in this task group o4-mini rejected 62% of tasks, which is higher than for the whole set (49%), hence the shorter mean difference between length of the ground truth and o4-mini solutions. Claude Sonnet 3.5, Gemini 2.5 Pro Preview, GPT-4o follow with the success rates 0.38-0.40. Six of the tested models have solutions on average more than twice as long as the reference solutions and the issues associated with this.
Table 5. Success rate in solving ’Spatial operations’ tasks
(53 tasks from total of 202) Model No Partial Match $\Delta$ in match match solution length Gemini 2.0 Flash 0.65 0.10 0.26 1.9 Gemini 2.5 Pro Preview 0.46 0.16 0.39 0.7 Claude Haiku 3.5 0.56 0.18 0.26 2.4 Claude Sonnet 3.5 0.50 0.11 0.40 1.6 Claude Sonnet 4 0.54 0.12 0.34 3.8 GPT-4o 0.55 0.06 0.38 2.3 GPT-4.1 0.46 0.19 0.35 2.9 o4-mini 0.33 0.10 0.56 -0.08
In the fourth task group, ’Spatial Operations’, the accuracy (success rate) for most of models is the lowest across all task groups (Table 6). The top three models in this task group are: o4-mini with an accuracy rate of 0.60, due to the better ability to identify unsolvable tasks, Gemini 2.5 Pro Preview, and GPT-4o.
@@
Table 6. Success rate in solving ’Heatmaps, contour lines’ tasks
(54 tasks from total of 202) Model No Partial Match $\Delta$ in match match solution length Gemini 2.0 Flash 0.74 0.05 0.21 1.3 Gemini 2.5 Pro Preview 0.60 0.08 0.31 0.8 Claude Haiku 3.5 0.72 0.13 0.15 2.0 Claude Sonnet 3.5 0.73 0.07 0.20 1.4 Claude Sonnet 4 0.65 0.15 0.20 4.6 GPT-4o 0.63 0.07 0.30 1.2 GPT-4.1 0.61 0.13 0.25 2.8 o4-mini 0.32 0.09 0.60 0.5
Table 7. Share of solutions matching across models and tasks groups
Model Unsolvable Solvable Merge - Process - Spatial Heatmaps,
tasks tasks visualize merge - operations contour
visualize lines
Gemini 2.0 Flash 0.40 0.39 0.68 0.48 0.26 0.21
Gemini 2.5 Pro Preview 0.55 0.39 0.67 0.50 0.39 0.31
Claude Haiku 3.5 0.22 0.40 0.60 0.36 0.26 0.15
Claude Sonnet 3.5 0.47 0.47 0.72 0.61 0.40 0.20
Claude Sonnet 4 0.17 0.55 0.71 0.41 0.34 0.20
GPT-4o 0.62 0.38 0.60 0.61 0.38 0.30
GPT-4.1 0.43 0.42 0.60 0.51 0.35 0.25
o4-mini 0.90 0.51 0.77 0.75 0.56 0.60
6.4. Token usage
We also compared the number of tokens used by the models to solve the benchmark set (Table 8) since it influences costs and costs are a critical parameter in practice. Note that because solving the tasks is an iterative process and at every iteration the previous model responses are fed back into model, the number of generated output tokens correlates with measured input tokens. What is clear is that an increase in token usage and generation does not necessarily translate into accuracy in solving tasks.
Anthropic models were on the higher end of the input/output tokens used. Claude Sonnet 4, due to its reasoning ability, creates 2-3 times more input tokens and up to 5 times more output tokens than competitors, yet achieves similar or worse accuracy than models like Claude Sonnet 3.5, GPT-4o, GPT-4.1 and Gemini 2.5 Pro Preview (Table 7). Anthropic models are more expensive in use: the most affordable of the three, Claude 3.5 Haiku, due to the amount of tokens generated ended up costing more than any OpenAI or Google models tested in this experiment to solve the benchmark set.
Notably, the best performed on this benchmark set OpenAI’s o4-mini model was among the three models with the lowest amount of input tokens used and the highest amount output token (3 times of the Claude Sonnet 4 output and 8 times GPT-41, next OpenAI’s model by output).
Table 8. Tokens used for generation of solutions for the whole benchmark set (single run)
@@ 6.4. Token usage
Gemini 2.0 Flash 7.5 40.7
Gemini 2.5 Pro Preview 10.4 85.4
Claude Haiku 3.5 14.9 196.0
Claude Sonnet 3.5 12.2 151.2
Claude Sonnet 4 26.9 291.6
GPT-4o 8.0 61.2
GPT-4.1 12.7 121.3
o4-mini 7.9 956.6
6.5. Error analysis
@@
  1. (1)
    Poor understanding of geometrical relations: e.g., taking values from centroid point of a polygon instead of calculating average value over polygon, trying to use centroids of lines while looking for geometries overlaps, not understanding that object’s centroid can lie outside the boundary of the object, etc.
  2. (2)
    Lack of recent world knowledge: LLMs worked well with established international groupings, but struggled with ambiguous groupings (Upper/Lower Nile countries) or with more recent changes (e.g., BRICS expansion in 2024), even if the event happened before the model’s cut-off date.
  3. (3)
    Over-reliance on the world knowledge: using internally known spelling of names rather than dataset values for filtering, despite having tools to check column values.
  4. (4)
    Inefficient data manipulations: merging without prefiltering (e.g., merging with all USA counties instead of first filtering to specific state); filtering by countries when filtering by continent/region was possible, and/or repeating filtering steps sometimes confusing LLM and derailing solutions.
  5. (5)
    Skipping steps that would narrow down the geodatasets to the required geography. E.g keeping all the North American lakes when the question was about USA.
  6. (6)
    Attempting to solve with partial data, trying to be helpful by using whatever geography was available instead of rejecting tasks with insufficient data coverage.
  7. (7)
    Modifying tasks to make them solvable rather than solving the original problem.
  8. (8)
    Hallucinating unjustified data processing steps, rejecting solvable tasks in the middle of solution, using some numbers for thresholds or names of countries in filters without justification or connection with provided datasets.
  9. (9)
    Answering/rejecting the task without using the tools and dataset provided.
Looking at models’ mistakes, it is likely that all models will perform much better if there is human-in-the-loop to disambiguate vague situations or provide step-by-step instructions on how to solve the task and more detailed explanations of the geographical scope of the tasks. It was also observed that making the tools more narrow specialized improved the accuracy of the agent.
7. Conclusions